OpenStudy (anonymous):

Hey Can Someone please HELP ME with these 4-t=3(t-1)-5 A 3 B 6 C NO SOLUTION D IDENTIFY -6A+3=3(2A-1) A a=5 B a=11 C no solution D identify

OpenStudy (ranga):

I will do the first one and you follow the same steps to do the second one. 4 - t = 3(t-1) - 5 4 - t = 3t - 3 - 5 4 - t = 3t - 8 4 + 8 = 3t + t 12 = 4t t = 3 So answer is A) 3

OpenStudy (anonymous):

thanks alot that really helped out , i was struggling lol

OpenStudy (anonymous):

hey SwoleFreshNation

OpenStudy (ilikecookies):

did you get it now